# MOR-VIT: Efficient Vision Transformer with Mixture-of-Recursions

### v1

@@ Abstract
Vision Transformers (ViTs) have achieved remarkable success in image recognition, yet standard ViT architectures are hampered by substantial parameter redundancy and high computational cost, limiting their practical deployment. While recent efforts on efficient ViTs primarily focus on static model compression or token-level sparsification, they remain constrained by fixed computational depth for all tokens. In this work, we present MoR-ViT, a novel vision transformer framework that, for the first time, incorporates a token-level dynamic recursion mechanism inspired by the Mixture-of-Recursions (MoR) paradigm. This approach enables each token to adaptively determine its processing depth, yielding a flexible and input-dependent allocation of computational resources. Extensive experiments on ImageNet-1K and transfer benchmarks demonstrate that MoR-ViT not only achieves state-of-the-art accuracy with up to 70% parameter reduction and 2.5× inference acceleration, but also outperforms leading efficient ViT baselines such as DynamicViT and TinyViT under comparable conditions. These results establish dynamic recursion as an effective strategy for efficient vision transformers and open new avenues for scalable and deployable deep learning models in real-world scenarios.
@@ 1 Introduction
To address this gap, we propose MoR-ViT, a novel Vision Transformer architecture that, for the first time, incorporates a token-level dynamic recursion mechanism inspired by the Mixture-of-Recursions (MoR) paradigm. In MoR-ViT, each token is routed through the network with a flexible, input-dependent recursive depth determined by a lightweight router, enabling adaptive allocation of computation based on token semantics. Our experiments on ImageNet-1K and transfer learning benchmarks show that MoR-ViT achieves competitive or superior accuracy to state-of-the-art efficient ViTs—including DynamicViT, TinyViT, and others—while delivering up to 70% parameter reduction and 2.5$\times$ inference acceleration. These results highlight dynamic recursion as a new and effective strategy for scalable, resource-efficient visual modeling, opening promising avenues for future research and deployment.
@@ 3.2 Mixture-of-Recursions (MoR) Mechanism
[…]
@@ 4.1 Experiment Setup
ViT-B/16 12 768 3072 12 $\times 16$ 16
MoR-B/16 (Ours) 12 768 3072 12 $\times 16$ 16
DeiT-B/16 12 768 3072 12 $\times 16$ 16
TinyViT-21M 32 384/512/768 1536/2048/3072 6-12 $\times 16$ 16
EfficientFormer-L1 12 384 1536 6 $\times 16$ 16
MetaFormer-S12 12 384 1536 6 $\times 16$ 16
TokenLearner-ViT-B 12 768 3072 12 $\times 16$ 16
@@ 4.2 Experiment Results
  • MoR-ViT (Ours)
  • ResNet-50 (ResNet)
  • Vision Transformer (ViT)
  • Data-efficient Image Transformer (DeiT)
  • DynamicViT
  • Adaptive Token Sampling (ATS)
  • EfficientFormer (EffFormer)
  • PoolFormer (MetaFormer)
  • MobileViT
  • TinyViT
  • TokenLearner
  • SparseFormer
  • Token Clustering Transformer (TCT)
@@
Model Params (M) FLOPs (G) img/s IN Top-1 (%) ADE20K mIoU (%)
@@ 5 Conclusion
Extensive experiments on ImageNet-1K and various downstream benchmarks demonstrate that MoR-ViT achieves state-of-the-art accuracy with up to 70% parameter reduction and 2.5× inference acceleration, outperforming leading efficient ViT baselines under comparable settings. Notably, these gains are achieved without resorting to external large-scale pretraining or distillation. Ablation studies confirm the critical contribution of the dynamic recursion mechanism and parameter sharing to the overall efficiency and performance.

### v2

@@ Abstract
Vision Transformers (ViTs) have achieved remarkable success in image recognition, yet standard ViT architectures are hampered by substantial parameter redundancy and high computational cost, limiting their practical deployment. While recent efforts on efficient ViTs primarily focus on static model compression or token-level sparsification, they remain constrained by fixed computational depth for all tokens. In this work, we present MoR-ViT, a novel vision transformer framework that, for the first time, incorporates a token-level dynamic recursion mechanism inspired by the Mixture-of-Recursions (MoR) paradigm. This approach enables each token to adaptively determine its processing depth, yielding a flexible and input-dependent allocation of computational resources. Extensive experiments on ImageNet-1K and transfer benchmarks demonstrate that MoR-ViT not only achieves state-of-the-art accuracy with up to 68% parameter reduction and 2.0× inference acceleration, but also outperforms leading efficient ViT baselines such as DynamicViT and TinyViT under comparable conditions. These results establish dynamic recursion as an effective strategy for efficient vision transformers and open new avenues for scalable and deployable deep learning models in real-world scenarios.
@@ 1 Introduction
[…]
@@ 3.2 Mixture-of-Recursions (MoR) Mechanism
MoR supports both expert-choice and token-choice routing schemes, with the former adopting hierarchical top-$k$ token selection at each recursion, and the latter assigning fixed recursion steps per token at the outset.
@@ 3.3 MoR-ViT Architecture
3.4 Implementation Considerations
This section details the key engineering decisions and architectural adaptations used to integrate the Mixture-of-Recursions (MoR) mechanism into a Vision Transformer (ViT) backbone. While §3.3 describes the core model, the items below specify routing, auxiliary objectives, recursion stability, masking, and sharing strategies that are critical to the robustness and efficiency of MoR-ViT.
3.4.1 Token Semantics and Routing Adaptation
Unlike language models where tokens denote words/subwords, ViTs operate on image patches. To make routing effective in the visual domain, we prepend a learnable [CLS] token as a global summary vector. This facilitates the router’s ability to detect spatially salient or semantically complex regions and to allocate deeper recursion to the corresponding tokens.
3.4.2 Auxiliary Losses and Routing Regularization
We use standard cross-entropy as the primary task loss and selectively incorporate auxiliary objectives inspired by the original MoR design:
z-loss: regularizes router logits to mitigate instability;
balancing loss: encourages a more even distribution of tokens across recursion depths;
sampling loss (optional): reduces route overfitting.
These are treated as regularizers with tunable weights. Although we have not observed pronounced routing imbalance in our vision experiments, we retain these mechanisms as preventive measures following best practice.
3.4.3 Stable Exit Strategy and Representation Caching
During recursive computation, a token’s hidden state is updated at each iteration. Deeper recursion can sometimes degrade representations (e.g., redundancy or attenuated attention). To avoid exit decisions being made on degraded vectors, we cache each token’s hidden state before the router at every step. When a token exits, the cached (pre-router) state is used as its final output rather than the potentially degraded post-block state.
3.4.4 Parameter Sharing with a Middle-Cycle Strategy
Instead of stacking many unique transformer blocks, MoR-ViT reuses a small set of shared blocks recursively (the “middle-cycle” strategy). This approach, inherited from MoR, reduces parameter count without sacrificing representational depth and contributes to both compression and efficiency.
3.4.5 Router Architecture and Activation Choice
We adopt a linear router with sigmoid activation, as recommended in the original MoR formulation. This design is lightweight and stable, avoiding the additional complexity of deep MLPs or softmax gating. Empirically, it performs reliably in vision tasks.
3.4.6 Recursive Exit Control and Token-Level Masking
To prevent exited tokens from interfering with subsequent attention, we employ token-level attention masks:
Exited tokens are removed from later block computations and from KV caches;
Computation thus focuses on active tokens only, saving FLOPs and memory.
3.4.7 Output Semantics and Loss Alignment
All tokens ultimately contribute to the loss, so their output vectors must remain semantically meaningful. By aligning the exit policy with the representation-caching mechanism, we ensure tokens exiting at different recursion depths still yield valid gradients, improving training stability.
@@ 4.1 Experiment Setup
ViT-B/16 12 768 3072 12 $16\times 16$
MoR-B/16 (Ours) 12 768 3072 12 $16\times 16$
DeiT-B/16 12 768 3072 12 $16\times 16$
TinyViT-21M 32 384/512/768 1536/2048/3072 6-12 $16\times 16$
EfficientFormer-L1 12 384 1536 6 $16\times 16$
MetaFormer-S12 12 384 1536 6 $16\times 16$
TokenLearner-ViT-B 12 768 3072 12 $16\times 16$
@@ 4.2 Experiment Results
MoR-ViT (Ours), ResNet-50 (ResNet), Vision Transformer (ViT), Data-efficient Image Transformer (DeiT), DynamicViT, Adaptive Token Sampling (ATS), EfficientFormer (EffFormer), PoolFormer (MetaFormer), MobileViT, TinyViT, TokenLearner, SparseFormer, Token Clustering Transformer (TCT).
@@
Model Params (M) FLOPs (G) img/s IN Top-1 (%) ADE mIoU (%)
@@ 4.3.5 Ablation Study: Effect of Dynamic Recursion and Parameter Sharing
We note that further experiments, including more extensive ablation studies and deployment evaluations, are planned for future work.
5 Discussion
In this section, we analyze the practical implications, limitations, and broader applicability of MoR-ViT, addressing both empirical observations and potential concerns.
5.1 Effectiveness of Dynamic Routing in Vision Tasks.
Our token-depth heatmaps reveal that regions with high texture complexity or salient semantic content are consistently assigned deeper recursion, while background or low-information regions tend to exit early. This token-wise alignment between recursion depth and feature difficulty confirms that dynamic routing is not only meaningful in natural language domains, but also well-suited for visual recognition tasks.
5.2 Performance Gains Beyond Parameter Scaling.
MoR-ViT achieves superior accuracy compared to the baseline ViT with fewer parameters. When parameters are matched, the performance advantage becomes even more pronounced. These results indicate that the gains originate from architectural efficiency—particularly the adaptive depth control—rather than from model scaling or parameter redundancy.
5.3 Inference Latency and Deployment Suitability.
The router is implemented as a lightweight linear module with minimal computational overhead. Most tokens exit after one or two recursion steps, and recursive blocks are parameter-shared. Consequently, the additional latency introduced by dynamic recursion is negligible. For deployment-critical scenarios, static thresholding and compilation-time graph unrolling can be employed to further optimize runtime performance.
5.4 Architectural Generalization to Other ViT Variants.
Because MoR-ViT modifies routing and recursion control without altering the underlying attention or feed-forward layers, it can be integrated into other ViT-family architectures, such as Swin, PVT, or TinyViT, with minimal adaptation. This design ensures broad applicability and facilitates cross-architecture adoption.
5.5 Applicability of KV Caching and KV Sharing.
Although KV caching and KV sharing are effective in the original MoR for language models, they are not directly beneficial in our visual setting. In ViT-based architectures, token dependencies and attention patterns differ substantially from autoregressive models, reducing the utility of KV reuse across recursion steps. For this reason, we omit these mechanisms in our implementation, prioritizing architectural simplicity and avoiding unnecessary complexity.
5.6 Representation Differences from Standard ViT Outputs.
While MoR-ViT produces output tensors with the same shape $(B,N,D)$ as standard ViTs, the semantic content is distinct. In standard ViTs, all tokens undergo identical processing through a fixed-depth stack, whereas MoR-ViT assigns adaptive recursion paths based on token complexity. As a result, the final representations encode heterogeneous computational histories, leading to richer and more differentiated token embeddings.
@@ 6 Conclusion
Extensive experiments on ImageNet-1K and various downstream benchmarks demonstrate that MoR-ViT achieves state-of-the-art accuracy with up to 68% parameter reduction and 2.0× inference acceleration, outperforming leading efficient ViT baselines under comparable settings. Notably, these gains are achieved without resorting to external large-scale pretraining or distillation. Ablation studies confirm the critical contribution of the dynamic recursion mechanism and parameter sharing to the strong efficiency and performance.